OpenStudy (lgbasallote):

Find the equation of the circle with center at (2,5) and tangent to the line 2x + 3y = 5

OpenStudy (eyust707):

Someone on here will be able to do this better than I but I can get you started

OpenStudy (eyust707):

we know the general form of a circle is \[(x-h)^2 + (y - k)^2 = r^2\]

OpenStudy (lgbasallote):

mmhmm

OpenStudy (eyust707):

where h and h are the x and y coordinates

OpenStudy (zepp):

 Like this?

OpenStudy (eyust707):

so therefore h=2 k = 5

OpenStudy (lgbasallote):

im thinking getting the slope of the tangent line then getting its perpendicular slope to get the slope of radius

OpenStudy (eyust707):

yes

OpenStudy (lgbasallote):

but how will i find the length then

OpenStudy (lgbasallote):

intersection?

OpenStudy (eyust707):

yes

OpenStudy (lgbasallote):

but how do i find the equation of the radius

OpenStudy (eyust707):

figure out a "b" that satisfies the center

OpenStudy (zepp):

\(2x + 3y-5 = 0\) (General form) -A/B to find the slope -2/3

OpenStudy (lgbasallote):

a "b" that satisfies the center?

OpenStudy (eyust707):

then use distance from center point to intesection

OpenStudy (eyust707):

lol there might be an easier way tho so stayed tuned

OpenStudy (lgbasallote):

hmm maybe i can use \[\frac{y_2 - y_1}{x_2 - x_1} = -2/3\]

OpenStudy (lgbasallote):

\[\frac{y_2 - 5}{x_2 - 2} = -2/3\]

OpenStudy (zepp):

The radius is perpendicular, so the sole of the radius is 3/2 Now find the equation of the radius, passing through (2,5) f(x) = mx+b 5=3/2(2)+b 5=3+b b=5-3=2 So equation of the radius is \(f(x)=\frac{3}{2}x+2\)

OpenStudy (lgbasallote):

lol i thought when you said slope is -a/b that was the perpendicular already -_-

OpenStudy (lgbasallote):

you should stop beating around the bush :P

OpenStudy (zepp):

Now find the point right there:  We have the tangent and the equation of the radius \(2x + 3y = 5\) and \(y=\frac{3}{2}x+2\) :P Sir lgba, http://openstudy.com/users/zepp#/updates/4fa3474fe4b029e9dc34125e